# Casper: Causality-Aware Spatiotemporal Graph Neural Networks for Spatiotemporal Time Series Imputation

### v1

@@ 3.2.3. Spatiotemporal Causal Attention
It can be theoretically proven that $\rho_{i^{\prime},t^{\prime}}$ will converge to 0 or 1 (see Section 3.3.1). Additionally, in practice, the correlation weight $\alpha$ can be easily extended to the multi-head version as in (Vaswani et al., 2017).
3.2.4. Loss Function
Casper is trained by the masked MAE:
@@ 3.2.4. Loss Function
where $m_{i,t}$ is the mask, $y_{i,t}$ is the ground-truth value, $\hat{y}_{i,t}$ is the predicted value, $\Phi$ is the set of all $\rho$ in the spatiotemporal causal attention, $\lambda$ is a tunable coefficient, and $||\cdot||_{1}$ denotes the $l_{1}$ norm.
@@ 3.3. Framework Analysis
3.3.1. Theoretical Analysis
We theoretically prove that $\rho_{i^{\prime},t^{\prime}}$ in Equation (16) will converge to 0 or 1 in Theorem 1. Therefore, $\rho_{i^{\prime},t^{\prime}}$ indicates the Granger causality. If $\rho_{i^{\prime},t^{\prime}}=0$, then $\beta_{i^{\prime},t^{\prime}}\cdot\alpha_{i^{\prime},t^{\prime}}=0$, showing that $\mathbf{h}^{in}_{i^{\prime},t^{\prime}}$ does not Granger cause $\mathbf{h}^{out}_{i,t}$. Moreover, from the proof of Theorem 1, it can be observed that $\rho_{i^{\prime},t^{\prime}}$ is actually a gradient based explanation (see Remark 1), which determines causal and non-causal relationships based on the gradients. Compared with the classic gradient based explanation methods (Selvaraju et al., 2017), which needs extra steps to calculate gradients after the model is trained, the proposed $\rho$ has two advantages: (1) $\rho$ does not require extra steps for calculating derivatives, and the value of $\rho$ could directly provide the explanation; (2) the parameters associated with $\rho$ are jointly trained with the model, and thus it can guide the model to focus on the most important relationships during training.
@@ 3.3.1. Theoretical Analysis
Due to space limitation, we only provide a proof sketch here. For simplicity, let’s only consider the loss for a single point $(i,t)$, where $m_{i,t}=1$ and ${y}_{i,t}-\hat{y}_{i,t}>0$:
@@
In gradient decent, the updating function of $w_{j}$ is:
@@
where $k$ is the iteration index and $\eta>0$ is the learning rate. As we only consider the parameters $\mathbf{w}$ of $\rho$, for simplicity, let’s fix all other parameters in the model. Therefore, ${\partial\hat{y}_{i,t}}/{\partial\mathbf{h}_{i,t}^{out}},\mathbf{v}_{i^{\prime},t^{\prime}},\alpha_{i^{\prime},t^{\prime}}$ are fixed. The normalization factor $Z\leq 1$ and $Z=1$ if and only if $\beta_{i^{\prime},t^{\prime}}=1$ for $\forall i^{\prime},t^{\prime}$. $\rho_{i^{\prime},t^{\prime}}(1-\rho_{i^{\prime},t^{\prime}})>0$ since $\rho_{i^{\prime},t^{\prime}}\in(0,1)$. Now, suppose $g>\lambda$, if $h_{j}^{in}>0$, then ${w}^{(k+1)}_{j}>{w}^{(k)}_{j}$. When $k\rightarrow\infty$, $w_{j}^{(k)}\rightarrow+\infty$. Otherwise, if $h_{j}^{in}<0$, then ${w}^{(k+1)}_{j}<{w}^{(k)}_{j}$. When $k\rightarrow\infty$, $w_{j}^{(k)}\rightarrow-\infty$. Therefore, when $k\rightarrow\infty$, $\mathbf{w}^{T}\mathbf{h}\rightarrow+\infty$. As a result, $\rho_{i^{\prime},t^{\prime}}^{(k)}=\sigma({\mathbf{w}^{(k)}}^{T}\mathbf{h}^{in})\rightarrow 1$, as $k\rightarrow\infty$. Similarly, if $g<\lambda$, then $\rho_{i^{\prime},t^{\prime}}^{(k)}$ will converge to 00. ∎
@@
3.3.2. Complexity Analysis
The complexities of the input project, the skip project and the transformer layer are $O(NT)$, $O(NT)$ and $O(NT^{2})$. The complexities of SCA and PBD are $O(ET)$ and $O(NN_{P}T)$, where $E$ is the number of edges in the sensor network, $N_{P}$ is the number of prompts. The overall complexity is $O(\max(E,NT,NN_{P})T)$. In SCA, if we calculate attention weights and causal gates for each pair of the data point without using $\mathbf{A}$, then the complexity will be extremely high: $O(N^{2}T^{2})\gg O(ET)$.
@@ 4.1. Experimental Setup
Datasets. Three public real-world datasets are used to evaluate Casper. AQI (Zheng et al., 2015) is an hourly record of air pollutants from 437 air quality monitoring stations in China from May 2014 to April 2015. Besides, we also use the popular AQI-36 (Cao et al., 2018) which is a reduced version of AQI containing records from 36 sensor stations scatter around Beijing. METR-LA (Li et al., 2018) contains traffic speed time series collected from 207 sensors on highways in Los Angeles for 4 months. PEMS-BAY (Li et al., 2018) is traffic speed time series collected from 325 sensors on highways in San Francisco Bay Area for 6 months. The records in both METR-LA and PEMS-BAY are collected every 5 minutes. For AQI, METR-LA and PEMS-BAY, the temporal window is set as $T=24$, and for AQI-36, the temporal window is 36. Following (Marisca et al., 2022), the sensor network is built based on the geographical distances between sensor stations.
@@
For AQI and AQI-36, we use the evaluation masks in (Yi et al., 2016; Marisca et al., 2022) which simulates the real missing data distribution in the datasets. We refer to this setting as the general missing. For METR-LA and PEMS-BAY, we consider both point missing and block missing settings as in (Marisca et al., 2022). In point missing, 25% data points are masked out. In block missing, 5% spatial blocks and 0.15% temporal blocks ranging from 1 hour to 4 hours are masked out.
@@
Evaluation Metrics. We use the standard Mean Absolute Error (MAE) and Mean Squared Error (MSE) between the ground truth values and the imputed values as the evaluation metrics.
@@
Baselines. We consider three groups of baselines. (1) Traditional statistical methods: the mean value of the sequence (MEAN); neighbor mean (KNN); matrix factorization (MF); multiple imputation using chained equations (MICE (White et al., 2011)); vector auto-regression (VAR); (2) Early deep learning models: rGAIN (Cini et al., 2021): an adversarial method similar to (Miao et al., 2021; Luo et al., 2018); BRITS (Cao et al., 2018): a bidirectional RNN method; (3) Recent deep learning models: ST-Transformer (Marisca et al., 2022): a spatiotemporal extension of the original Transformer (Vaswani et al., 2017); GRIN (Cini et al., 2021): a graph enhanced recurrent neural network; SPIN (Marisca et al., 2022): a spatiotemporal graph attention based imputation model. Whenever possible, the results of baselines are copied from the corresponding paper.
@@
Implementation Details. We set the size of embeddings as 32. The numbers of layers for the encoder for AQI-36 and other datasets are 2 and 4 respectively. Following (Marisca et al., 2022), we use the Adam optimizer (Kingma and Ba, 2014) with a learning rate of 0.0008 and a cosine scheduler to train the model. The maximum number of epochs is 300 and the patience of early stopping is 40 epochs. Batch size is fixed as 8. During training, $p\in[0.2,0.5,0.8]$ data points are randomly masked out for each batch, and the loss is calculated based on these masked points. Our code is implemented based on PyG (Fey and Lenssen, 2019) and Torch Spatiotemporal (Cini and Marisca, 2022). Code will be released upon publication.
@@ 4.2. Overall Performance
The overall performance (MAE and MSE) of different methods is presented in Table 1. The upper, middle and lower groups of baselines are the traditional statistical methods, the RNN methods and the recent methods (Transformer based and graph based methods). Generally speaking, the RNN methods perform better than the statistical methods, and the recent methods further outperform the RNN methods. When imputing a data point, these methods exploit all the available information in the context, without identifying the causal relationships between the data point and the context. However, it is inevitable that some confounders are included in the data, such as noise and non-causal shortcut edges. Over-reliance on the confounders could lead to overfitting and make the model susceptible to noise. The proposed Casper could effectively remove the impact of confounders. As shown in Table 1, Casper achieves the lowest overall MAE and MSE scores and also has lower standard deviations, demonstrating the effectiveness of enforcing the model to discover causality during training.
@@
4.3. Ablation Study
Effectiveness of PBD,SCA and $\mathbf{A}$. In the upper part of Table 2, we investigate the impact of the Prompt-Based Decoder (PBD), the Spatiotemporal Causal Attention (SCA) and the sensor network $\mathbf{A}$. First, significant performance drop on MAE and MSE can be observed when we remove PBD (replace PBD with a MLP) and/or SCA (remove the causal gate $\beta$) from the full model Casper, indicating the effectiveness of the frontdoor adjustment and the causal gate for improving the overall performance. It is also worth noting that, on MSE, the standard deviations of the ablated versions of Casper (w/o PBD and/or SCA) are significantly higher than the full model Casper, demonstrating that enforcing the model to focus on the causality could improve its robustness. Second, although $\mathbf{A}$ contains non-causal relationships among sensors, it still has critical contributions to the overall performance, demonstrating the necessity of considering $\mathbf{A}$ when imputing missing values.
Effectiveness of the Prompts. In the lower part of Table 2, we demonstrate the effectiveness of using the learnable prompts to capture the global contextual information of the datasets by replacing the prompts with other approximations, including K-means cluster centers and randomly sampled data. To obtain the other two approximations, we first pre-train an imputation model, i.e., Casper without SCA and PBD, and then extract the embedding of each training sample $\mathbf{G}$ by applying average/max pooling over the embeddings of all the points in $\mathbf{G}$. For the cluster center approximation, we apply K-means over the emebddings to obtain 1,000 cluster centers. For the sampling approximation, we randomly sample 1,000 embeddings for each training sample. Compared with the prompts, the cluster centers and randomly sampled embeddings not only perform worse but also require extra efforts to obtain embeddings, which demonstrates the superiority of prompts.
@@ 4.4. Visualization
In this subsection, we provide visualization to further analyze Casper’s ability for discovering the causality.
@@
Discovered Causal Relationships. In Figure 6, we visualize the discovered causal relationships corresponding to Figure 4(b) and 4(f). The AQI-36 dataset is collected from 36 air quality monitoring stations in Beijing, which records PM2.5 levels. We draw the sensors on the map of Beijing according to their latitudes and longitudes. The widths of the arrows correspond to the values of attention scores. It is evident that for the query sensors 14 and 1 in the two figures, not every their nearby sensor has a causal relationship with them. In Figure 5(a), although the sensor 6, 5, 15 are very close to the query 14, Casper discovers that there is no causal relationship among them. We conjecture that this is because the wind blew generally westward and northward in Beijing during the time period corresponding to the input data. In Figure 5(b), the neighboring sensors 24, 15, 6, 20, 9, 13 are not regarded as the causes for the query sensor 1. According to the spatial relationships, the information of the sensors 15, 6, 20, 9, 13 might be included in the causal sensors such as 2, 5 14, 8, since the causal sensors are in between with the query and theses non-causal sensors. However, there is no other sensor between 1 and 24. By taking a closer look at the upper-left corner of the map, we find that the sensors 1 and 24 are separated by the Fragrant Hills. Therefore, the air quality at the sensor 1 might not be directly influenced by the sensor 24 for that the time period corresponding to the input data. The two examples in Figure 6 show that the proposed Casper could effectively discover the causal relationships among sensors, which provides better insights for further data analysis. Additionally, the two examples also show that the simple distance based sensor network construction is biased, which contains many non-causal correlations, since it ignores other factors in the real world, such as wind direction and terrain.

### v3

@@ 3.2. Architecture of Casper
It can be theoretically proven that $\rho_{i^{\prime},t^{\prime}}$ will converge to 0 or 1 (see Section 3.3). Additionally, in practice, the correlation weight $\alpha$ can be easily extended to the multi-head version as in (Vaswani et al., 2017).
Loss Function
Casper is trained by the masked MAE. For a given spatiotemporal time series segment with $N$ nodes and $T$ length, the loss is defined as:
@@
where $m_{i,t}$ is the mask, $y_{i,t}$ is the ground-truth value, $\hat{y}_{i,t}$ is the predicted value, $\Phi$ is the set of all $\rho$ in SCA, $\lambda$ is a tunable coefficient, and $||\cdot||_{1}$ denotes the $l_{1}$ norm.
@@ 3.3. Framework Analysis
Theoretical Analysis
We theoretically prove that $\rho_{i^{\prime},t^{\prime}}$ in Equation (16) will converge to 0 or 1 in Theorem 1, and thus $\rho_{i^{\prime},t^{\prime}}$ indicates the Granger causality. If $\rho_{i^{\prime},t^{\prime}}=0$, then $\beta_{i^{\prime},t^{\prime}}\cdot\alpha_{i^{\prime},t^{\prime}}=0$, showing that $\mathbf{h}^{in}_{i^{\prime},t^{\prime}}$ does not Granger cause $\mathbf{h}^{out}_{i,t}$. Moreover, from the proof of Theorem 1, it can be observed that $\rho_{i^{\prime},t^{\prime}}$ is actually a gradient-based explanation (see Remark 1), which determines causal and non-causal relationships based on the gradients. Compared with the classic gradient-based explanation methods (Selvaraju et al., 2017), which needs extra steps to calculate gradients after the model is trained, the proposed $\rho$ has two advantages: (1) $\rho$ does not require extra steps for calculating derivatives, and the value of $\rho$ could directly provide the explanation; (2) the parameters associated with $\rho$ are jointly trained with the model, and thus it can guide the model to focus on the most important relationships during training.
@@
For simplicity, let’s only consider the loss for a single point $(i,t)$, where $m_{i,t}=1$ and ${y}_{i,t}-\hat{y}_{i,t}>0$:
@@
where $\mathbf{h}_{i,t}^{out},\beta_{i^{\prime},t^{\prime}}$ are from Equation (17). In gradient descent, the updating function of $w_{j}$ is:
@@
where $k$ is the iteration index and $\eta>0$ is the learning rate. As we only consider the parameters $\mathbf{w}$ of $\rho$, for simplicity, let’s fix all other parameters in the model. Therefore, ${\partial\hat{y}_{i,t}}/{\partial\mathbf{h}_{i,t}^{out}},\mathbf{v}_{i^{\prime},t^{\prime}},\alpha_{i^{\prime},t^{\prime}}$ are fixed. The normalization factor $Z\leq 1$, and $Z=1$ if and only if $\beta_{i^{\prime},t^{\prime}}=1$ for $\forall i^{\prime},t^{\prime}$. $\rho_{i^{\prime},t^{\prime}}(1-\rho_{i^{\prime},t^{\prime}})>0$ since $\rho_{i^{\prime},t^{\prime}}\in(0,1)$. Now, suppose $g>\lambda$, if $h_{j}^{in}>0$, then ${w}^{(k+1)}_{j}>{w}^{(k)}_{j}$. When $k\rightarrow\infty$, $w_{j}^{(k)}\rightarrow+\infty$. Otherwise, if $h_{j}^{in}<0$, then ${w}^{(k+1)}_{j}<{w}^{(k)}_{j}$. When $k\rightarrow\infty$, $w_{j}^{(k)}\rightarrow-\infty$. Therefore, when $k\rightarrow\infty$, $\mathbf{w}^{T}\mathbf{h}\rightarrow+\infty$. As a result, $\rho_{i^{\prime},t^{\prime}}^{(k)}=\sigma({\mathbf{w}^{(k)}}^{T}\mathbf{h}^{in})\rightarrow 1$, as $k\rightarrow\infty$. Similarly, if $g<\lambda$, then $\rho_{i^{\prime},t^{\prime}}^{(k)}$ will converge to 00. ∎
@@
Complexity Analysis
The complexities of the input project, the skip project, and the transformer layer are $O(NT)$, $O(NT)$, and $O(NT^{2})$. The complexities of SCA and PBD are $O(ET)$ and $O(NN_{P}T)$, where $E$ is the number of edges in the sensor network, and $N_{P}$ is the number of prompts. The overall complexity is $O(\max(E,NT,NN_{P})T)$. In SCA, if we calculate attention weights and causal gates for each pair of the data points without using $\mathbf{A}$, then the complexity will be extremely high: $O(N^{2}T^{2})\gg O(ET)$.
@@ 4.1. Experimental Setup
In this subsection, we briefly explain the datasets, evaluation metrics, and baselines used for the experiments.
Datasets.
@@
Three public real-world benchmark datasets are used to evaluate Casper. AQI (Zheng et al., 2015) is an hourly record of air pollutants from 437 air quality monitoring stations in China from May 2014 to April 2015. We also use the popular AQI-36 (Cao et al., 2018) which is a reduced version of the full AQI containing records from 36 sensor stations scattered around Beijing. METR-LA (Li et al., 2018) contains traffic speed time series collected from 207 sensors on highways in Los Angeles for 4 months. PEMS-BAY (Li et al., 2018) is a traffic speed time series collected from 325 sensors on highways in San Francisco Bay Area for 6 months. The time series records in METR-LA and PEMS-BAY are collected every 5 minutes. For AQI, METR-LA, and PEMS-BAY, the temporal window is set as $T=24$, and for AQI-36, the temporal window is 36. To be consistent with prior works (Marisca et al., 2022), the adjacency matrices of sensor networks are built by applying a thresholded Gaussian kernel (Shuman et al., 2013; Li et al., 2018) over the geographical distances between sensor stations.
@@
For AQI and AQI-36, we use the evaluation masks in (Yi et al., 2016; Marisca et al., 2022) which simulates the real missing data distribution in the datasets. We refer to this setting as the general missing. For METR-LA and PEMS-BAY, we consider both points missing and block missing settings as in (Marisca et al., 2022). In point missing, 25% data points are masked out. In block missing, 5% spatial blocks and 0.15% temporal blocks ranging from 1 hour to 4 hours are masked out.
Evaluation Metrics.
@@
We use the standard Mean Absolute Error (MAE) and Mean Squared Error (MSE) between the ground truth values and the imputed values as the evaluation metrics.
Baselines.
@@
We consider three groups of methods as our baselines. (1) Traditional statistical methods: the mean value of the sequence (MEAN); neighbor mean (KNN); matrix factorization (MF); multiple imputations using chained equations (MICE (White et al., 2011)); vector auto-regression (VAR); (2) Early deep learning models: rGAIN (Cini et al., 2021): an adversarial method similar to (Miao et al., 2021; Luo et al., 2018); BRITS (Cao et al., 2018): a bidirectional RNN imputation method; (3) Recent deep learning models: ST-Transformer (Marisca et al., 2022): a spatiotemporal extension of the original Transformer (Vaswani et al., 2017); GRIN (Cini et al., 2021): a graph enhanced recurrent neural network; SPIN (Marisca et al., 2022): a spatiotemporal graph attention based imputation model; PoGeVon (Wang et al., 2023): a recent spatiotemporal imputation method which is based on the position-aware spatiotemporal graph variational auto-encoder. Whenever possible, the results of baselines are copied from the corresponding paper.
Implementation Details.
@@
Most of the training configurations follow prior works (Marisca et al., 2022). We set the size of embeddings as 32. The numbers of layers for the encoder for AQI-36 and other datasets are 2 and 4 respectively. We use the Adam optimizer (Kingma and Ba, 2014) with a learning rate of 0.0008 and a cosine scheduler to train the model. The maximum number of epochs is 300 and the patience of early stopping is 40 epochs. Batch size is fixed as 8. During training, $p\in[0.2,0.5,0.8]$ data points are randomly masked out for each batch, and the loss is calculated based on these masked points.
@@ 4.2. Overall Performance
In this subsection, we show the overall performance (MSE and MAE) of Casper for the imputation task to demonstrate the overall competence of Casper for imputation.
The overall performance of different methods is presented in Table 1. The upper, middle, and lower groups of baselines are the traditional statistical methods, the RNN methods and the recent methods (Transformer based and graph based methods). Generally speaking, the RNN methods perform better than the statistical methods, and the recent methods further outperform the RNN methods. When imputing a data point, these methods exploit all the available information in the context, without identifying the causal relationships between the data point and the context. However, it is inevitable that some confounders are included in the data, such as the non-causal shortcut edges. Over-reliance on the confounders could lead to overfitting and make the model susceptible to noise. The proposed Casper could effectively remove the impact of confounders. As shown in Table 1, Casper achieves the lowest overall MAE and MSE scores and also has lower standard deviations, demonstrating the effectiveness of enforcing the model to discover causality during training.
@@
[…]
w/o skip project 10.35$\pm$0.21 438.02$\pm$28.38
4.3. Ablation Study
Effectiveness of PBD,SCA and $\mathbf{A}$.
In the upper part of Table 2, we investigate the impact of the Prompt-Based Decoder (PBD), the Spatiotemporal Causal Attention (SCA) and the sensor network $\mathbf{A}$. First, a significant performance drop on MAE and MSE can be observed when we remove PBD (replace PBD with an MLP) and/or SCA (remove the causal gate $\beta$) from the full model Casper, indicating the effectiveness of the frontdoor adjustment and the causal gate for improving the overall performance. It is also worth noting that, on MSE, the standard deviations of the ablated versions of Casper (w/o PBD and/or SCA) are significantly higher than the full model Casper, demonstrating that enforcing the model to focus on the causality could improve its robustness. Second, although $\mathbf{A}$ contains non-causal relationships among sensors, it still has critical contributions to the overall performance, demonstrating the necessity of considering $\mathbf{A}$ when imputing missing values.
Effectiveness of the Prompts.
In the middle part of Table 2, we demonstrate the effectiveness of using the learnable prompts to capture the global contextual information of the datasets by replacing the prompts with other approximations, including K-means cluster centers and randomly sampled data. To obtain the other two approximations, we first pre-train an imputation model, i.e., Casper without SCA and PBD, and then extract the embedding of each training sample $\mathbf{G}$ by applying average/max pooling over the embeddings of all the points in $\mathbf{G}$. For the cluster center approximation, we apply K-means over the embeddings to obtain 1,000 cluster centers. For the sampling approximation, we randomly sample 1,000 embeddings for each training sample. Compared with the prompts, the cluster centers and randomly sampled embeddings not only perform worse but also require extra effort to obtain embeddings, which demonstrates the superiority of prompts.
Effectiveness of Other Components.
In the lower part of Table 2, we study the impact of time constraints and the skip project layer in the model. First, if we enforce the time constraint, i.e., $t^{\prime}\leq t$ in Definition 3.1, for the imputation task, then the performance will significantly drop. These results demonstrate that it is vital to take into consideration both the past and future points for the imputation task. Second, removing the skip project layer will have negative impacts on the overall performance, showing the power of the skip project, which aligns with the observation of the residue connections in the literature (He et al., 2016; Jing et al., 2021c; Yan et al., 2021a).
@@ 4.4. Visualization
In this subsection, we provide visualization to further analyze Casper’s ability of causality discovery.
Attention Maps.
@@
We draw the most salient causal relationships in Figure 5(b) and 5(f) on the map of Beijing in Figure 6(a) and 6(b). Each number corresponds to a sensor and the arrow width corresponds to the attention weights. It is evident that for the query sensors 14 and 1 in the two figures, not every nearby sensor has a causal relationship with them. In Figure 6(a), although the sensors 6, 5, 15 are very close to the query 14, Casper discovers that there is no causal relationship among them. We conjecture that this is because the wind blew generally westward and northward in Beijing for Figure 5(b) and 5(f) (Zhang et al., 2012). In Figure 6(b), the neighboring sensors 24, 15, 6, 20, 9, 13 are not regarded as the causes for the query sensor 1. According to the spatial relationships, the information of the sensors 15, 6, 20, 9, 13 might be included in the causal sensors such as 2, 5 14, and 8, since the causal sensors are in between with the query and these non-causal sensors. However, there is no other sensor between 1 and 24. By taking a closer look at the upper-left corner of the map, we find that sensors 1 and 24 are separated by the Fragrant Hills. Therefore, the air quality at sensor 1 might not be directly influenced by sensor 24 for the period of the input data. The two examples in Figure 6 show that the proposed Casper could effectively discover the causal relationships among sensors, which provides better insights for further data analysis. Additionally, the two examples also show that the simple distance-based sensor network construction is biased, which contains many non-causal correlations since it ignores other factors in the real world, such as wind direction and terrain.
4.5. Causal Graph Discovery on the Quasi-Realistic Dataset
In the real world, the ground truth causal graphs are usually unavailable, and thus it is difficult to quantitatively evaluate the ability of causal graph discovery. Therefore, we follow (Cheng et al., 2022) and evaluate our Casper on a quasi-realistic data called DREAM-3 (Prill et al., 2010), which is a gene expression data regulation dataset.
DREAM-3 contains $N=100$ gene expression levels and the length of the expressions is $T=21$. The goal is to discover the causal relationship among the 100 gene expression levels. For DREAM-3, we train Casper via the imputation task, and use similar training configurations as AQI-36. Following (Cheng et al., 2022), we use AUC between the ground-truth graph and the discovered graph as the evaluation metric. Specifically, for Casper we obtain a causal weight matrix $\mathbf{C}_{i,t}\in\mathbb{R}^{N\times T}$ for each data point $x_{i,t}$, where $\mathbf{C}_{i,t}[i^{\prime},t^{\prime}]=\beta_{i^{\prime},t^{\prime}}$ in Equation 17. By concatenating all the $\mathbf{C}_{i,t}$ together, where $i\in\{1,\cdots,N\}$ and $t\in\{1,\cdots,T\}$, we have a tensor $\mathcal{C}\in\mathbb{R}^{N\times T\times N\times T}$, where the first two dimensions correspond to the data point $x_{i,t}$, and the last two dimensions correspond to the causal weight $\mathbf{C}_{i,t}$ for $x_{i,t}$. We obtain the final causal weight matrix by max pooling over the two dimensions corresponding to the time, and normalizing by $T^{2}$: $\mathbf{A}=\text{max-pool}(\mathcal{C},\text{dim=}2,4)/T^{2}\in\mathbb{R}^{N\times N}$.
The results of Casper and several SOTA time series causal discovery baselines are presented in Table 3, where the results of baselines are copied from (Cheng et al., 2022). Table 3 shows that Casper achieves the highest AUC score. This experiment quantitatively demonstrates that Casper has a strong ability of discovering causal relationships.
Table 3. Causal Graph Discovery on DREAM-3.
Models PCMCI(Runge et al., 2019) NGC(Tank et al., 2021) eSRU(Khanna and Tan, 2019) LCCM(De Brouwer et al., 2020) NGM(Bellot et al., 2021) CUTS(Cheng et al., 2022) Casper
AUC 0.5517 0.5579 0.5587 0.5046 0.5477 0.5915 0.6325
4.6. More Results
In this subsection, we provide more experimental results of Casper, including convergence of $\rho$, and sensitivity analysis.
Convergence of $\rho$.
We set $0.1$ and $0.9$ as the thresholds to round $\rho$. Specifically, if $\rho\leq 0.1$ then we regard $\rho$ has converged to 00; similarly, if $\rho\geq 0.9$, then we regard $\rho$ has converged to $1$. The statistical results of the AQI-36 dataset show that 98% $\rho$ converges to 0 or 1, which corroborates Theorem 1.
Sensitivity Analysis.
We present the results of sensitivity experiments of $\lambda$ and the number of prompts $N_{P}$ in Figure 7(a)-7(b). For $\lambda$, the lowest MAE can be obtained when $\lambda$ is around 0.001. For $N_{P}$, the best results can be obtained when $N_{P}\in[200,1400]$.